four equal charge, Q = 2 microcoulomb,are located at the corners of a square that is 0.50m on a side. What is the potential energy of the array?


1
Expert's answer
2022-01-14T09:46:13-0500

The potential energy of this configuration is


"U=2·\dfrac 1{4\pi\epsilon_0}\dfrac{q^2}{a}+\dfrac 1{4\pi\epsilon_0}\dfrac{q^2}{a\sqrt2}=0.19\text{ J}."